# Machine-Learning Driven Load Shedding to Mitigate Instability Attacks in Power Grids

### v1

@@ Abstract
Every year critical infrastructure becomes more complex and we grow to rely on it more and more. With this reliance, it becomes an attractive target for cyberattacks from sophisticated actors, with one of the most attractive targets being the power grid. One class of attacks, instability attacks, is a newer type of attack that has relatively few protections developed. We present a cost effective, data-driven approach to training a supervised machine learning model to retrofit load shedding decision systems in power grids with the capacity to defend against instability attacks. We show a proof of concept on the IEEE 14 Bus System using the Achilles Heel Technologies Power Grid Analyzer, and show through an implementation of modified Prony analysis (MPA) that MPA is a viable method for detecting instability attacks and triggering defense mechanisms.
@@ I Introduction
With this potential disaster and its consequences in mind, it is a troubling idea that wide scale power outages could be purposefully brought about by terrorists or nation-state sponsored attackers, and a yet further troubling idea given that attacks like these have already happened at large scales [4].
Because of this, it is critical to come to understand what kinds of attacks may be around the corner in the ever growing landscape of cyber-warfare targeting critical infrastructure, and that is what this paper sets out to do.
@@
This paper presents a novel retrofitting of current remedial action mechanisms, namely load shedding mechanisms, using a cost-effective, data-driven approach to be able to respond to attacks before they do significant damage and interrupt them with 92% overall accuracy. Methods for attack detection, as well as tools for creating training data of the attacks, are presented, which allow operators to train models on their own systems and establish software based logic to be able to respond to this new class of attack.
@@ II Related Work
A much more extensive catalog exists regarding using load shedding with machine learning, the approach taken in this paper. Schweitzer et al. [12], in conjunction with their development of modified Prony analysis, use load shedding to combat unstable inter-area oscillations. This is seen in other works as well; Toro et al. implemented automatic load shedding for the development of more robust remedial action schemes [13], which gives precedence for using load shedding without an human operator in the loop to combat growing oscillations. Justin and Paternain [14] implement a reinforcement learning (RL) agent in conjunction with a classifier of system response given a certain input to do adaptive load shedding. This approach is very similar to the work presented here, and integrating the classifier here with a similar type of RL agent could prove useful in broad scale implementation with the hereafter mentioned ’operator-type’ entity. The principal difference between their paper and this one is that we focus on training a classifier on instability attacks (with infrastructure to detect these attacks)—system responses that would almost certainly not be captured in traditional system configurations used in that paper. There are many other things happening in this space, a systematic survey of which was done recently by Skrjanc et al.[15].
@@
Figure 1: (a) Top: An example of naively shedding according to under voltage load shedding schemes under an instability attack. $y$ here is the voltage of the third bus in the IEEE 14 Bus System. The tenth load is shed at 10 seconds, and the attack worsens relative to what it was before. (b) Bottom: An example of a stable load shed with the same system initial conditions and topology as (a) under a different attack, shedding the same load, showing the same value in the $y$ axis as in (a). The behavior of (a) and (b) are reflected in all the other signals in the other buses, generators, and so on, to differing scalings. [^2] For angular values such as $\delta$, they enter into typical behaviors post load shedding that require things like governor dynamics to return to stability, behaviors not captured in this model
Figure 2: The IEEE 14 Bus system, modeled in the AHT Power Grid Analyzer tool. Red circles correspond to generators, white nodes refer to buses, each with a load.
@@ III Methodology and Threat Model
As a high-level anecdote to explain the tool, imagine that an adversary establishes the groundwork for an instability attack and launches it. Alarms are triggered by the modified Prony analysis mechanisms that are normally used for operational grid stability but have had the relay logic adjusted slightly to detect these advanced attacks. Data begins to be captured and is fed into a machine learning model (either cloud or edge based) that quickly provides an operator-type entity (be it an actual operator or another decision-making algorithm) a list of loads to shed that would stop the attack, and which loads (if shed) would worsen the attack. The entity then makes a decision, shedding a load, and the attack excites into a transient that passes, yielding way to normal post-load shedding procedures and steady-state stability.
This is one way the tool presented in this paper could be used and demonstrates the different parts at play here. We have an alarm system that uses modified Prony analysis to detect an attack and trigger a machine learning model. That model takes in system states for a given amount of time and yields a labeling for each load: stable shed, or unstable shed (as in, if we shed this load, will the system stabilize or get worse). That can then be used by an operator or a decision-making algorithm outside the scope of this paper to make a decision informed by all factors and stop an instability attack.
@@ III-A Threat Model
We therefore consider a zero-trust environment, where we assume that at least two nodes in our system have been compromised and we do not know which ones. Our action is to shed loads (via flipping breakers), and we assume our capacity to do so has largely not been compromised. This assumption is somewhat fair, as the attacker likely must use the same infrastructure to carry out the attack that a defender would use to stop the attack, and so it likely cannot be completely disabled, and even if it is, we consider load shedding opportunities across the power grid, and compromising them at scale is also unlikely.
@@ III-B Motivating Example
As a motivating example, consider the IEEE 14 Bus System shown in Figure 2, a popular research benchmark in the power systems community for studying a variety of phenomena, including stability. Physically, shedding a load is a matter of a breaker being tripped, ceasing the draw of power for a given community on the grid. In our simulation of the IEEE 14 Bus System, this amounts to setting the load states on a given bus to zero. In traditional load shedding schemes, loads to be shed are chosen a priori to any event, and are almost entirely dependent on the topology of the system. If we take a similar approach and shed the same bus in the same topology with the same initial conditions, the traditional approach suggests that we would get the same outcome. But the reality, as seen in Figure 2, is that for two different instability attacks, shedding the same bus gives different results: one outcome where the system returns to stability, and the other where the attack is amplified. This shows that choosing a load to shed under an instability attack is a non-trivial problem, and needs to be addressed. One way to address this is through a data-driven approach.
@@ III-C Data Source and Processing
For a data-driven approach, we need data. To show a proof-of-concept tool that demonstrates the power and efficacy of data-driven loadshedding to bypass instability attacks, we used the Achilles Heel Technologies (AHT) Power Grid Analyzer, a tool that can take in a model of power grid and analyze its vulnerability to instability attacks, including modeling the results of these attacks. I modified this tool to be able to implement load shedding by cutting to zero at a discrete time in the numerical solver the values for a given load. This is meant to simulate the tripping of a breaker, the principal way load shedding is done in the real world. While no model is a substitute for reality, this numerical simulation of load shedding under an instability attack yields valuable insight and promise for further developed tools to be based on this approach. As a standard preset for all of our modeling, we chose the IEEE 14 Bus System with presets given by the software tool. More about the tool can be found in Tackett et al. [9].
@@
For the sake of training our model, there is essentially three pieces of information we start off with: 1) which load was shed, 2) the system states for the attack (before and after load shedding) and 3) the type of attack. The data we would eventually want to feed into the model for training then consisted of three somewhat separate parts: 1) which load would be shed, 2) the system states before the load shed, and 3) whether or not the load shed would lead to a stable outcome [^3] Preliminary surveying the parameter space found that at stages of the attack when the signal was still small, the response to load shedding was independent of time, so ten seconds after the attack was launched was used as the universal load shedding time for data generation and model training.. Ultimately then, going from the data we had to the data we needed was trivial for the first two components of the model training data: we would iterate through every type of attack for a given system with every load shedding possibility, collect all of them, and chop the data off for the pre-load shedding point. However, an issue remains: how would we know how to trigger our model?
III-C1 Modified Prony Analysis (MPA)
@@
A method used for finding unstable inter-area oscillations proved to be able to establish an ‘alarm’ of sorts to alert a system that an attack is underway. This method is called the modified Prony analysis (MPA), presented by the Schweitzer Engineering Lab (SEL) [12]. The paper presents a method for a data-driven method (not machine learning) that establishes whether or not a given signal has an unstable mode. Prony analysis was originally used in ringdown analysis, where an impulse would be put into a system, and the modes would then be analyzed. It does this by doing a parametric curve fitting that leverages that any signal can be approximated by damped complex exponentials. Modified Prony analysis does a similar thing, but instead of using an impulse to analyze stable modes, it looks at a sliding window of the data and identifies unstable modes early on. It does this in an iterative way that rejects noise modes according to some tolerances and hyperparameters, as described in the SEL paper [12].
@@
III-C2 Heuristic Label
The heuristic label [^4] The heuristic label used here is specifically designed for the quirks and limitations of the numerical integrator and edge case handling of the AHT Power Grid Analyzer and are presented for clarity alone. Using this heuristic for any other system is not recommended and it is instead recommended to look into using MPA as described in the previous section. That said, if a reader would like all the code used for the generation of data and the heuristic that is then applied to get the training labels, such code is available upon request to the author for reproducibility’s sake, with exceptions and corresponding concessions applying to where the AHT Power Grid Analyzer source code is proprietary. can be divided up into three tests:
@@ III-C2 Heuristic Label
III-C3 Data Processing
With the heuristic label in hand and now having all the tools to generate the data we want, we created a multi-threaded script to process all of the data and put it into training, testing, and validation data sets for the machine learning model.
@@ III-C3 Data Processing
First, if the integrator stopped after the load shedding for any reason (numerical stiffness being a principal culprit), that wouldn’t work because there would be no way to label the data. Further to this aspect, if the data didn’t go on for at least 50 seconds after the load shed (most would go to about 200), then we decided not to use it as labeling would likely be inconclusive. Finally, if the integrator just ran into problems for some reason, we decided to see how much data we could successfully get and if it was sufficient to not need to address the bugs that caused those to fail.
@@ V-A Model Results
After training the model for several hours on a Macbook Pro with an M3 Pro chip and 18 gigabytes of memory, the model was able to correctly classify loads 92% of time accurately, as shown in detail in Table I.
@@ V-B Data Statistics
Next, we can see that while we could only consider about half of the total attack/ load configurations, every read and write variable was able to be represented in some way. It is probable that those lower in representation were affected by issues that caused many of them to be cut, like mentioned in Section III-C3, and this could also affect the cases where there is a significantly lower amount of successful load sheds present, but also may be indicative of some attacks just being harder to defend against with single load shedding than others. This is a topic of great interest, and will be explored in future work.
@@ VI Discussion
The primary contributions of this paper are the model infrastructure in code (which can be found in the GitHub link cited on the first page), the proof of concept of this model on the IEEE 14 Bus System using the AHT Power Grid Analyzer, and tools for creating data for other systems (available on request for reasons mentioned above). A secondary contribution that serves more as an extension of research done by others is the identification of the modified Prony analysis as an effective alarm system and data labeler to train and trigger a classifier on early stage attack data. These tools do some things well, while leaving other areas for future research.
@@ VI-A Strengths
The model does well in a couple regards that we had mentioned before: by using a data-driven approach, we sidestep computational time requirements, model variance issues (inasmuch as the system trained on is allowed to vary), and complexity issues by utilizing supervised machine learning to create a labeler based on system states.
@@ VII Conclusion
While instability attacks are not prevalent today, in a world of critical infrastructure increasingly becoming a target in cyberwarfare, their discovery and implementation by powers that wish to do harm is inevitable, so finding ways to mitigate against them is critical. We have presented a data-driven model that can use load shedding systems in a power grid to identify how to respond against these attacks, laying down the foundation and proving the concept for industry level tools to apply these approaches at scale and identify ways to prevent these attacks from doing serious damage to critical infrastructure around the world. We have shown that naive load shedding can be disastrous, and given tools for detecting these attacks as well as labeling their affects post load shedding, as well as simulating these attacks in testbeds to get the highest quality data for training more extensive models. We have identified strengths and weaknesses, and identified ways these systems must be improved for the practical problem of implementing the solution at scale.

### v2

@@ Abstract
Critical infrastructures are becoming increasingly complex as our society becomes increasingly dependent on them. This complexity opens the door to new possibilities for attacks and a need for new defense strategies. Our work focuses on instability attacks on the power grid, wherein an attacker causes cascading outages by introducing unstable dynamics into the system. When stress is place on the power grid, a standard mitigation approach is load-shedding: the system operator chooses a set of loads to shut off until the situation is resolved. While this technique is standard, there is no systematic approach to choosing which loads will stop an instability attack. This paper addresses this problem using a data-driven methodology for load shedding decisions. We show a proof of concept on the IEEE 14 Bus System using the Achilles Heel Technologies Power Grid Analyzer, and show through an implementation of modified Prony analysis (MPA) that MPA is a viable method for detecting instability attacks and triggering defense mechanisms.
@@ I Introduction
With this potential disaster and its consequences in mind, it is a troubling idea that wide scale power outages could be purposefully brought about by terrorists or nation-state sponsored attackers, and a yet further troubling idea given that attacks like these have already happened at large scales [4, 5].
Because of this, it is critical to come to understand what kinds of attacks may be around the corner in the ever growing landscape of cyber-warfare targeting critical infrastructure, and how such attacks might be mitigated.
@@
This paper presents a novel retrofitting of current remedial action mechanisms, namely load shedding mechanisms, using a cost-effective, data-driven approach to be able to respond to attacks before they do significant damage and interrupt them with a 0.92 overall F1-score. Methods for attack detection are also presented. These allow operators to train models on their own systems and establish software based logic to be able to respond to this new class of attack.
@@ II Related Work
A much more extensive catalog exists regarding using load shedding with machine-learning, the approach taken in this paper. Schweitzer et al. [13], in conjunction with their development of modified Prony analysis, use load shedding to combat unstable inter-area oscillations. This is seen in other works as well; Toro et al. implemented automatic load shedding for the development of more robust remedial action schemes [14], which gives precedence for using load shedding without an human operator in the loop to combat growing oscillations. Justin and Paternain [15] implement a reinforcement learning (RL) agent in conjunction with a classifier of system response given a certain input to do adaptive load shedding. This approach is very similar to the work presented here, and integrating the classifier here with a similar type of RL agent could prove useful in broad scale implementation with the hereafter mentioned ’operator-type’ entity. The principal difference between their paper and this one is that we focus on training a classifier on instability attacks (with infrastructure to detect these attacks)—system responses that would almost certainly not be captured in traditional system configurations used in that paper. There are many other things happening in this space, a systematic survey of which was done recently by Skrjanc et al.[16].
@@
Figure 1: An example of naively shedding according to under voltage load shedding schemes under an instability attack. On the left is the IEEE 14 Bus system, with generators as smaller red circles and buses as the white nodes. Two attacks are shown; a red $\Delta$ and a yellow $\Delta$, where each $\Delta$ reads from a state in a given node (Bus 5 and Generator 3 for yellow and red $\Delta$’s, respectively) and both write to the same state in Bus 9; these are all highlighted accordingly. These attacks happen separately, with the same response: shedding the load on Bus 6. The resulting effect on the system, shown by the voltage for Bus 3 (in normalized units), is shown in the two graphs to the right. In the instance of the red $\Delta$, the attack gets worse, and in the instance of the yellow $\Delta$, the attack is interrupted, thereby showing that using pre-defined load-shedding schemes can yield stable or unstable outcomes, and a system-aware decision process is needed.
@@ III Methodology and Threat Model
As a high-level anecdote to explain the tool, imagine that an adversary establishes the groundwork for an instability attack and launches it. Alarms are triggered by the MPA mechanisms that are normally used for operational grid stability but have had the relay logic adjusted slightly to detect these advanced attacks. Data begins to be captured and is fed into a machine-learning model (either cloud or edge based) that quickly provides an operator-type entity (be it an actual operator or another decision-making algorithm), a list of loads to shed that would stop the attack, and which loads (if shed) would worsen the attack. The entity then makes a decision, shedding a load, and the attack excites into a transient that passes, yielding way to normal post-load shedding procedures and steady-state stability.
This is one way the tool presented in this paper could be used and demonstrates the different parts at play here. We have an alarm system that uses MPA to detect an attack and trigger a machine-learning model. That model takes in system states for a given amount of time and yields a labeling for each load: stable shed, or unstable shed (as in, if we shed this load, will the system stabilize or get worse). That can then be used by an operator or a decision-making algorithm outside the scope of this paper to make a decision informed by all factors and stop an instability attack.
@@ III-A Threat Model
We therefore consider a zero-trust environment, where we assume that at least two nodes in our system have been compromised and we do not know which ones. Our action is to shed loads (through physical actions like flipping breakers), and we assume our capacity to do so has largely not been compromised. This assumption is somewhat fair, as the attacker likely must use the same communication network and other infrastructure to carry out the attack that a defender would use to stop the attack, and so it likely cannot be completely disabled, and even if it is, we consider load shedding opportunities across the power grid, and compromising them at scale is a scenario not within the scope of this threat model.
@@ III-B Motivating Example
As a motivating example, consider the IEEE 14 Bus System shown in Figure 1, a popular research benchmark in the power systems community for studying a variety of phenomena, including stability. Physically, shedding a load is a matter of a breaker being tripped, ceasing the draw of power for a given community on the grid. In our simulation of the IEEE 14 Bus System (the dynamics for which can be found in Appendix A), this amounts to setting the load states (the active and reactive loads, referred to in this paper as $PL$ and $QL$) on a given bus to zero at a discrete point in time. In traditional load shedding schemes, loads to be shed are chosen a priori to any event, and are almost entirely dependent on the topology of the system. If we take a similar approach and shed the same bus in the same topology with the same initial conditions, the traditional approach suggests that we would get the same outcome. But the reality, as seen in Figure 1, is that for two different instability attacks, shedding the same bus gives different results: one outcome where the system returns to stability, and the other where the attack is amplified and even exacerbated. This shows that choosing a load to shed under an instability attack is a non-trivial problem, and needs to be addressed. One way to address this is through a data-driven approach.
@@ III-C Data Source and Processing
For a data-driven approach, we need data. To show a proof-of-concept tool that demonstrates the power and efficacy of data-driven loadshedding to bypass instability attacks, we used the Achilles Heel Technologies (AHT) Power Grid Analyzer, a tool that can take in a model of power grid and analyze its vulnerability to instability attacks, including modeling the results of these attacks. We modified this tool to implement load shedding by cutting to zero at a discrete time in the numerical solver the values for a given load. While no model is a substitute for reality, this numerical simulation of load shedding under an instability attack yields valuable insight and promise for further developed tools to be based on this approach. We used the generic, predefined generator models and parameter values given in the software tool to model the IEEE 14 Bus System. More about the tool can be found in Tackett et al. [10].
@@
For the sake of training our machine-learning model, there are essentially three pieces of information we start off with: 1) which load was shed, 2) the system states for the attack (before and after load shedding) and 3) the states being read or written to by $\Delta$. The data we would eventually want to feed into the model for training then consisted of three somewhat separate parts: 1) which load would be shed, 2) the system states before the load shed, and 3) whether or not the load shed would lead to a stable outcome. We obtained the first two by iterating through each type of attack and every load shedding possibility, simulating each such scenario, and truncating the data at a well-defined pre-load shedding point.
Modified Prony Analysis (MPA)
@@
Next, we considered how the process to trigger the model. A method used for finding unstable inter-area oscillations proved to be able to establish an ‘alarm’ of sorts to alert a system that an attack is underway. This method is called the modified Prony analysis (MPA), presented by the Schweitzer Engineering Lab (SEL) [13]. The paper presents a method for a data-driven method (not machine-learning) that establishes whether or not a given signal has an unstable mode. Prony analysis was originally used in ringdown analysis, where an impulse would be put into a system, and the modes would then be analyzed. It does this by doing a parametric curve fitting that leverages that any signal can be approximated by damped complex exponentials. Modified Prony analysis does a similar thing, but instead of using an impulse to analyze stable modes, it looks at a sliding window of the data and identifies unstable modes early on. It does this in an iterative way that rejects noise modes according to some tolerances and hyperparameters, as described in the SEL paper [13].
@@
Heuristic Label
The heuristic label [^2] The heuristic label used here is specifically designed for the quirks and limitations of the numerical integrator and edge case handling of the AHT Power Grid Analyzer and are presented for clarity alone. If a reader would like all the code used for the generation of data and the heuristic that is then applied to get the training labels, such code is available upon request to the author for reproducibility’s sake, with exceptions and corresponding concessions applying to where the AHT Power Grid Analyzer source code is proprietary. can be divided up into three tests:
@@
Data Processing
With the heuristic label in hand and now having all the tools to generate the data we want, we created a multi-threaded script to process all of the data and put it into training, testing, and validation data sets for the machine-learning model.
@@
First, if the integrator stopped after the load shedding for any reason (numerical stiffness being a principal culprit), that wouldn’t work because there would be no way to label the data. Further to this aspect, if the data didn’t go on for at least 50 seconds [^3] Preliminary surveying the parameter space found that at stages of the attack when the signal was still small, the response to load shedding was independent of time, so ten seconds after the attack was launched was used as the universal load shedding time for data generation and model training. after the load shed (most would go to about 200), then we decided not to use it as labeling would likely be inconclusive. Finally, if the integrator just ran into problems for some reason, we decided to see how much data we could successfully get and if it was sufficient to not need to address the bugs that caused those to fail.
@@ V-A Model Results
After training the model for several hours on a Macbook Pro with an M3 Pro chip and 18 gigabytes of memory, the model was able to correctly classify loads with an overall F1-score of 0.92, as shown in detail in Table I.
@@ V-B Data Statistics
Next, we can see that while we could only consider about half of the total attack/ load configurations, every read and write variable was able to be represented in some way. It is probable that those lower in representation were affected by issues that caused many of them to be cut, like mentioned in Section III-C, and this could also affect the cases where there is a significantly lower amount of successful load sheds present, but also may be indicative of some attacks just being harder to defend against with single load shedding than others. This is a topic of great interest, and will be explored in future work.
@@ VI Discussion
The primary contributions of this paper are the model infrastructure in code and the proof of concept of this model on the IEEE 14 Bus System using the AHT Power Grid Analyzer. A secondary contribution that serves more as an extension of research done by others is the identification of the MPA as an effective alarm system and data labeler to train and trigger a classifier on early stage attack data. These tools do some things well, while leaving other areas for future research.
@@ VI-A Strengths
The model does well in a couple regards that we had mentioned before: by using a data-driven approach, we sidestep computational time requirements, model variance issues (inasmuch as the system trained on is allowed to vary), and complexity issues by utilizing supervised machine-learning to create a labeler based on system states.
@@ VII Conclusion
While instability attacks are not prevalent today, in a world of critical infrastructure increasingly becoming a target in cyberwarfare, their discovery and implementation by powers that wish to do harm seems inevitable, so finding ways to mitigate against them is critical. We have presented a data-driven model that can use load shedding systems in a power grid to identify how to respond against these attacks, laying down the foundation and proving the concept for industry level tools to apply these approaches at scale and identify ways to prevent these attacks from doing serious damage to critical infrastructure around the world. We have shown that naive load shedding can be disastrous, and given tools for detecting these attacks as well as labeling their effects post load shedding. We have identified strengths and weaknesses, and identified ways these systems must be improved for the practical problem of implementing the solution at scale.
Appendix A Power Grid Dynamic System
Generator equations (with AVR):
$v_{d}$ $=V\sin(\delta-\theta)$ (1)
$v_{q}$ $=V\cos(\delta-\theta)$ (2)
$P_{g}$ $=v_{d}i_{d}+v_{q}i_{q}$ (3)
$Q_{g}$ $=v_{q}i_{d}-v_{d}i_{q}$ (4)
$\tau_{e}$ $=(r_{a}i_{q}+v_{q})i_{q}+(r_{a}i_{d}+v_{d})i_{d}$ (5)
$\dot{\delta}$ $=\Omega_{b}(\omega-\omega_{s})$ (6)
$\dot{\omega}$ $=\frac{1}{2H}(\tau_{m}-\tau_{e}-D(\omega-\omega_{s}))$ (7)
$\dot{e}^{\prime}_{q}$ $=\frac{1}{T^{\prime}_{d0}}(-e^{\prime}_{q}-(x_{d}-x^{\prime}_{d})i_{d}+v_{f})$ (8)
$\dot{e}^{\prime}_{d}$ $=\frac{1}{T^{\prime}_{q0}}(-e^{\prime}_{d}+(x_{q}-x^{\prime}_{q})i_{q})$ (9)
$0$ $=v_{q}+r_{a}i_{q}-e^{\prime}_{q}+x^{\prime}_{d}i_{d}$ (10)
$0$ $=v_{d}+r_{a}i_{d}-e^{\prime}_{d}-x^{\prime}_{q}i_{q}$ (11)
$\dot{v}_{m}$ $=\frac{1}{T_{m}}\left(V-v_{m}\right)$ (12)
$\dot{v}_{r1}$ $=\frac{1}{T_{a}}\left(K_{a}\left(v_{\text{ref}}-v_{m}-v_{r2}-\frac{K_{f}}{T_{f}}v_{f}\right)-v_{r1}\right)$ (13)
$\dot{v}_{r2}$ $=-\frac{1}{T_{f}}\left(\frac{K_{f}}{T_{f}}v_{f}+v_{r2}\right)$ (14)
[…]
$v_{r}$ $=\begin{cases}v_{r}^{\text{min}}&v_{r1}<v_{r}^{\text{min}}\\ v_{r1}&v_{r}^{\text{min}}<v_{r1}<v_{r}^{\text{max}}\\ v_{r}^{\text{max}}&v_{r1}>v_{r}^{\text{max}}.\end{cases}$ (16)
Bus equations ($i$ is bus index):
$0$ $=P_{g,i}-P_{\ell,i}+\sum_{k}P_{b,ik}$ (17)
$0$ $=Q_{g,i}-Q_{\ell,i}+\sum_{k}Q_{b,ik}.$ (18)
Transmission line equations:
$P_{b,ik}$ $=Y_{ik}V_{i}(V_{k}\cos(\phi_{ik}+\theta_{i}-\theta_{k})-V_{i}\cos(\phi_{ik}))$ (19)
$Q_{b,ik}$ $=Y_{ik}V_{i}(V_{k}\sin(\phi_{ik}+\theta_{i}-\theta_{k})-V_{i}\sin(\phi_{ik}))+b_{ik}V_{i}^{2}.$ (20)
Transformer sending side equations:
$P_{b,ik}$ $=Y_{ik}\frac{V_{i}}{m_{ik}}\left(V_{k}\cos(\phi_{ik}+\theta_{i}-\theta_{k})-\frac{V_{i}}{m_{ik}}\cos(\phi_{ik})\right)$ (21)
$Q_{b,ik}$ $=Y_{ik}\frac{V_{i}}{m_{ik}}\left(V_{k}\sin(\phi_{ik}+\theta_{i}-\theta_{k})-\frac{V_{i}}{m_{ik}}\sin(\phi_{ik})\right)+b_{ik}V_{i}^{2}.$ (22)
Transformer receiving side equations:
$P_{b,ik}$ $=Y_{ik}V_{i}\left(\frac{V_{k}}{m_{ik}}\cos(\phi_{ik}+\theta_{i}-\theta_{k})-V_{i}\cos(\phi_{ik})\right)$ (23)
$Q_{b,ik}$ $=Y_{ik}V_{i}\left(\frac{V_{k}}{m_{ik}}\sin(\phi_{ik}+\theta_{i}-\theta_{k})-V_{i}\sin(\phi_{ik})\right).$ (24)